# Hyperparameter Optimization in Machine Learning

### v2

@@ 4.5 Discussion
Model-based methods tend to be somewhat more complex than other HPO techniques, both conceptually and from an implementation standpoint. They require fitting a (probabilistic) surrogate model with relatively few function evaluations, and to update this fit online whenever new observations become available. Moreover, the optimal decision on where to evaluate $f$ next needs to be approximated via an AF depending on the surrogate model, as we saw in Section 4.2.3. Thus, while there is a thriving ecosystem of free academic software, it is not rare that model-based HPO techniques are implemented and proposed as commercial products [sigopt-web-page, golovin2017google, AMT2020]. The development of MB techniques requires several design choices to be made, such as the concrete representation of $\hat{f}$. Some configurations may be left to be set by the final user. While authors often provide “default values”, the presence of many such parameters may make some model-based techniques less accessible. See, e.g., the list of required inputs for the algorithms proposed by hazan2017hyperparameter or falkner2018bohb.
While the foremost representatives of these type of methods – tied to BO – are classically black-box procedures, model-based techniques may greatly benefit from direct access to the underlying learning algorithm. For instance, several works [swersky2014freeze, klein2017fast, falkner2018bohb] use information gathered from partial execution of $A$ to stop unpromising evaluations, which may be additionally resumed at a later stage [swersky2014freeze] as we will discuss in more details in the next chapter. Before doing so, we spend some words on what is possibly the biggest practical limitation of (naive) model-base methods: parallelizability.
@@ 4.5.1 Decision making under parallelism
Independent of how configurations are suggested, synchronous job scheduling tends to be substantially less efficient than asynchronous. Namely, the time to process a batch is determined by the slowest one, while up to $W-1$ workers are idle waiting. Such differences in training times are particularly pronounced when searching over parameters determining the model size. In an asynchronous system, whenever a worker becomes available, it can be assigned a new configuration. While BO needs to make single suggestions only (as opposed to batches), this has to be done in the presence of up to $W-1$ pending evaluations, i.e. jobs which have been started for previously suggested configurations, but did not return a metric value yet. Ignoring pending evaluations runs the same risk of selecting highly redundant configurations in subsequent rounds. Fortunately, dealing with pending evaluations is quite related to ensuring diversity in batch selection, and local penalty methods apply just as well [Gonzales:16, Alvi:19]. Simpler approaches start from the observation that in sequential BO, diversity is enforced by conditioning on observed points. The resulting very low posterior variance shrinks the AF to small values near such points. In the kriging believer heuristic [Ginsbourger:10, Desautels:14], we impute the response for each pending evaluation by its current posterior mean and condition on this pseudo-datapoint. A somewhat more principled approach is fantasizing [snoek_practical_2012, Wilson2018], where pending target values are integrated out by Monte Carlo: they are sampled several times from the joint posterior over pending configurations, and the AF is averaged over these samples.
@@
One of the main challenges of HPO is that training and validating a machine learning algorithm each time the objective function is evaluated can be expensive. This is particularly true for modern machine learning algorithms applied to large datasets, such as deep neural networks, which can often take hours or days to train. In this chapter, we consider how to obtain and exploit low-fidelity signals that serve as approximations of the objective function. These auxiliary metrics are much faster to evaluate than computing the full objective function and often provide a strong enough signal to guide the optimization process.
@@ 5.1 Multi-fidelity optimization
As noted in Section 2.5, an important aspect of manual search and hyperparameter tuning is fast turnaround time. To achieve this, practitioners typically work with a subset of the training data during their initial explorations. They also do not wait for training to fully converge but instead check validation performance after a few iterations or epochs. While these early metrics are not entirely reliable, they are useful for identifying promising regions of the search space.
In this section, we discuss how these manual practices can be transformed into algorithmic tools that enable automated early stopping of underperforming runs.
@@
While we can freely choose $b$ during the optimization process, our primary interest lies in the performance at the full budget $b_{max}$. More formally, our goal is (still) to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$, which coincides with the formalization of Chapter 2.
In the next section, we explore commonly used fidelity measures that can be leveraged to accelerate the optimization process.
@@ 5.1.1 Iterations
For iterative machine learning algorithms, such as neural networks, we can interpret the number of iterations as the budget and use the performance at each iteration as an approximation of the final performance [swersky2014freeze, li2017hyperband, jamieson2016non]. In the case of neural networks, we typically consider epochs as iterations rather than individual mini-batch updates to reduce the overhead of validation error computations.
@@ 5.1.2 Training dataset subsets
The runtime of most machine learning algorithms scales linearly or even super-linearly with the number of training data points. Figure 5.1 (right) shows the validation error of a support vector machine trained on increasing subsets of the MNIST dataset, with each color representing a different hyperparameter configuration. As expected, the validation error decreases as the training dataset size increases. More interestingly, however, the ranking of hyperparameter configurations remains relatively consistent across dataset subsets [bornschein-icml20].
In the context of deep neural networks, this observation is related to recent research on scaling laws [kaplan-arxiv19, hoffmann-arxiv22], which empirically demonstrate that the training loss of autoregressive transformer models [vaswani_attention_2017] often follows a power-law function with respect to the size of the training dataset.
@@ 5.2.1 Successive halving
=$\mathcal{R}=\{b_{min}\eta^{k}\mid k\in[K]\}$
We assume, for simplicity, that $b_{max}/b_{min}=\eta^{K}$, where $K\in\mathbb{N}$. Here, $\eta\geq 2$ defines the halving constant and determines how aggressively SH prunes configurations.
@@
1:Initial budget $b_{min}$, maximum budget $b_{max}$
2:Set of $n$ candidate configurations $\mathcal{C}_{0}=\{\lambda_{1},\lambda_{2},\dots\lambda_{n}\}$
3:Halving constant $\eta\geq 2$
4:for $i\in[|\mathcal{R}|]$ do
5:     $\{f_{r_{i}}(\lambda):\lambda\in\mathcal{C}_{i-1}\}$triangleright Evaluate configurations on the rung
6:     $k\leftarrow\lfloor|\mathcal{C}_{i-1}|/\eta)\rfloor$triangleright Number of configurations to keep
7:     $\mathcal{C}_{i}\leftarrow\underset{\lambda\in\mathcal{C}_{i-1}}{\arg\mathrm{top}\text{-}k}f_{r_{i}}(\lambda)$triangleright Keep $k$ best performing configurations
8:end for
Algorithm 2 shows pseudo code for SH. Given a set $\mathcal{C}_{0}$ of $n$ hyperparameter configurations, usually drawn uniformly at random, we evaluate them at the first rung level, i.e., each configuration is evaluated with a budget of $b_{min}$ (Line 2). For convenience, the number of initial configurations is often set to $n=\eta^{K}$, but this can also be arbitrary. Afterwards, SH promotes $\frac{1}{\eta}$ of the configurations, sorted by their validation performance, to the next higher rung level $b_{min}\eta$ (Line 4). These steps are iterated until we reach the highest rung level, $b_{max}$. See Figure 5.2 for a visualization.
While conceptually simple, SH has shown strong performance in the literature and even outperformed more sophisticated Bayesian optimization techniques that only operate on the full budget [jamieson2016non, li2017hyperband]. However, its performance relies on a high correlation between the ranking of hyperparameter configurations across rung levels. In other words, the performance of a configuration at the first rung level should be representative of its performance at the highest rung level. Otherwise, configurations that perform best at $b_{max}$ may be discarded too quickly.
Figure 5.2: Successive halving splits the maximum budget $b_{max}$ into a set of exponentially increasing rungs. Then, in each iteration, successive halving evaluates all configurations on the current rung and promotes the top $\frac{1}{\eta}$ of the configurations to the next rung.
@@ 5.2.2 Hyperband
The minimum budget $b_{min}$ allocated to each configuration is an important meta-parameter of SH. If $b_{min}$ is too small, we may filter out good configurations based on unreliably low fidelities, while if $b_{min}$ is too large, our advantage over standard random search may diminish. Hyperband [li2017hyperband] addresses this issue by running different instances, $l=0,\dots,K$, of SH in sequence. The $l$-th instance (or bracket) has rung levels
$\mathcal{R}_{l}=\{b_{min}\eta^{k}\mid k\in\{l,\dots,K\}\},$
so $l=0$ corresponds to SH with $K+1$ rung levels, while $l=K$ corresponds to random search with a single level $b_{max}$. Each bracket has its own number of initial configurations $n_{l}$, which is chosen such that the total budget spent in each bracket is roughly the same. The risk of choosing too small a value for $b_{min}$ is mitigated at a modest extra expense.
@@ 5.2.3 Synchronous successive halving
How is parallel computation exploited by SH? Recall the notation of synchronous versus asynchronous parallelization from Section 2.4, and assume that $W$ workers are available. For synchronous scheduling, we can start $\min(W,M)$ jobs in parallel, where $M$ is the number of free slots remaining in the currently processed rung. Here, $W-M$ workers remain idle, which is particularly wasteful for the higher rungs. For example, if $n=\eta^{K}$, the highest rung has only a single slot, so $W-1$ workers will be idle for the maximum budget $b_{max}$. The same issue arises for asynchronous scheduling. A remedy will be discussed in Section 5.2.4.
Hyperband inherits the idle worker issue noted above for synchronous scheduling. However, since brackets are independent of each other, we can work on several of them in parallel, which (at the expense of some implementation complexity) eliminates idle time for asynchronous scheduling.
@@ 5.2.4 Asynchronous successive halving (ASHA)
In ASHA, all available computational resources are utilized immediately, ensuring that workers are never idle. Additionally, superior configurations are quickly promoted toward $b_{max}$ without the need to wait for synchronization points. In practice, ASHA often significantly outperforms synchronous SH or Hyperband [Klein:20].
@@ 5.3.1 Fabolas
Compared to classical Bayesian optimization (see Chapter 4), we cannot simply optimize standard acquisition functions as described in Section 4.2.3. For example, in the case of expected improvement, the highest improvements are expected at $b=b_{max}$. As a result, this approach would lead to over-exploitation, primarily evaluating configurations at the highest budget.
@@
$p_{\min}(\lambda^{*}\mid\mathcal{C})=\mathbb{P}(\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f(\lambda))$
@@ 5.3.2 BO-HB: Bayesian Optimization Hyperband
More specifically, once the proposed method BO-HB observes $N_{\min}$ data points at the rung level $b_{i}$, it trains two kernel density estimators—$l(\lambda)$ and $g(\lambda)$—to model the probability of a configuration $\lambda$ achieving a validation error lower or higher than a certain threshold, as defined in Eq. 4.16. However, the density estimators are only trained on all data points at the given rung level.
@@
While the original implementation of TPE by bergstra2011algorithms used a univariate kernel density estimator, falkner2018bohb proposed using a multivariate kernel density estimator to better capture interactions between hyperparameters. To fit such a model, at least $N_{\min}=m+1$ data points are required, where $m$ is the dimensionality of the search space.
@@
In this chapter we discuss a class of randomized methods for global optimization that evolve a population of candidate solution, drawing inspiration from the process of natural evolution [back1997handbook, simon2013evolutionary]. Evolutionary algorithms date back to the beginning of computer science [neumann1966theory] and have been used to tackle non-smooth and non-convex problems across many domains. Methods in this class typically rely on function evaluations only. Their application to hyperparameter optimization includes methods like CMA-ES [hansen1996adapting], population-based training (PBT) [jaderberg2017population] and particle swarm optimization [lorenzo2017particle], which we review below.
@@ 6.1 Evolutionary algorithms
Traditionally, evolutionary algorithm (EA) distinguish between evolutionary strategies (ES) and genetic algorithms (GA). GA refers to a class of methods that deal with binary problems and discrete objective functions while ES is concerned with real-valued, continuous objectives.
The terminology of EA techniques is borrowed from biology: the set of points is called population and configurations are called individuals, chromosomes or genotypes. Iterations are called generations and the objective function is called fitness. Typically at the end of each generation the fitness of all individuals is observed, making EA inherently parallel algorithms – traditional implementations usually requiring synchronization. After observing the fitness of the population, a series of transformations is applied to the individuals. Common types of transformations, which collectively implement explore/exploit strategies in this context, include:
@@
Other strategies like elitism (keep the best individuals unchanged) and niching (maintain a diverse population) are often used in conjunction to the explore/exploit strategies outlined above.
We described this class of algorithm by the procedure outlined in Population-Based-HPO in Algorithm 3. W can think about individuals and generations as draws from $N$ distributions $p_{k}^{j}$, where $N$ is the population size. These distributions implement the strategies described above, and evolve at each generation using global information from the current generation. For instance, elitism can be represented by setting $p_{k}^{j}=\delta_{\lambda_{k}^{j}}$ for the best performing configurations, where $\delta$ is a Dirac delta distribution. In contrast to random or quasi-random search methods covered in Chapter 3, population-based methods keep in general $N$ distinct distributions (rather than a global one) that are updated iteratively (rather than fixed).
@@
6:         $\lambda_{k}^{j}\sim p^{j}_{k-1}$triangleright Draw an individual
7:         Evaluate $f(\lambda_{k}^{j})$triangleright Evaluate its fitness (if needed)
@@
9:     Update $p^{j}_{k}$ from $p^{j}_{k-1}$ using $\{(\lambda^{i}_{k},f(\lambda^{i}_{k}))\}_{i\in[N]}$triangleright “Evolve” distributions
@@
Various implementations of ES/GA have been proposed in HPO e.g. for tuning hyperparameters of support vector machines [friedrichs2005evolutionary], for neural architecture search [real2019regularized, miikkulainen2024evolving] (see Section 8.3 for a discussion on neural architecture search), parameters of games [kunanusont2017ntuple] and even entire machine learning algorithms [real2020automl]. In the next sections, we review three particular implementations that have received notable attention.

### v3

@@ 4.4 Discussion
Model-based methods tend to be somewhat more complex than other HPO techniques, both conceptually and from an implementation standpoint. They require fitting a (probabilistic) surrogate model with relatively few function evaluations, and to update this fit online whenever new observations become available. Moreover, the optimal decision on where to evaluate $f$ next needs to be approximated via an AF depending on the surrogate model, as we saw in Section 4.2.3. Thus, while there is a thriving ecosystem of free academic software, it is not rare that model-based HPO techniques are implemented and proposed as commercial products [sigopt-web-page, golovin2017google, AMT2020]. The development of model-based techniques requires several design choices to be made, such as the concrete representation of $\hat{f}$. Some configurations may be left to be set by the final user. While authors often provide “default values”, the presence of many such parameters may make some model-based techniques less accessible. See, e.g., the list of required inputs for the algorithms proposed by hazan2017hyperparameter or falkner2018bohb.
While the foremost representatives of these type of methods — tied to BO — are classically black-box procedures, model-based techniques may greatly benefit from direct access to the underlying learning algorithm. For instance, several works [swersky2014freeze, klein2017fast, falkner2018bohb] use information gathered from partial execution of $A$ to stop unpromising evaluations, which may be additionally resumed at a later stage [swersky2014freeze] as we will discuss in more details in the next chapter. Before doing so, we discuss what is possibly the biggest practical limitation of conventional model-base methods: parallelizability.
@@ 4.4.1 Decision Making Under Parallelism
Independent of how configurations are suggested, synchronous job scheduling tends to be substantially less efficient than asynchronous. Namely, the time to process a batch is determined by the slowest one, while up to W1W-1 workers are idle waiting. Such differences in training times are particularly pronounced when searching over parameters determining the model size. In an asynchronous system, whenever a worker becomes available, it can be assigned a new configuration. While BO needs to make single suggestions only (as opposed to batches), this has to be done in the presence of up to W1W-1 pending evaluations, i.e. jobs which have been started for previously suggested configurations, but did not return a metric value yet. Ignoring pending evaluations runs the same risk of selecting highly redundant configurations in subsequent rounds. Fortunately, dealing with pending evaluations is quite related to ensuring diversity in batch selection, and local penalty methods apply just as well [Gonzales:16, Alvi:19]. Simpler approaches start from the observation that in sequential BO, diversity is enforced by conditioning on observed points. The resulting very low posterior variance shrinks the AF to small values near such points. In the kriging believer heuristic [Ginsbourger:10, Desautels:14], we impute the response for each pending evaluation by its current posterior mean and condition on this pseudo-datapoint. A somewhat more principled approach is fantasizing [snoek_practical_2012, Wilson2018], where pending target values are integrated out by Monte Carlo: they are sampled several times from the joint posterior over pending configurations, and the AF is averaged over these samples.
@@
One of the main challenges of HPO is that training and validating a machine learning algorithm each time the objective function is evaluated can be expensive. This is particularly true for modern machine learning algorithms applied to large datasets, such as deep neural networks, which can often take hours or days to train. In this chapter, we consider how to obtain and exploit low-fidelity signals that serve as approximations of the objective function. These auxiliary metrics are faster to evaluate and often provide a strong enough signal to guide the optimization process.
@@ 5.1 Multi-Fidelity Optimization
As noted in Section 2.5, an important aspect of manual search and hyperparameter tuning is a fast turnaround time. To achieve this, practitioners typically work with a subset of the training data during their initial explorations. They also do not wait for training to fully converge but instead check validation performance after a few iterations or epochs. While these early metrics are not entirely reliable, they are useful for identifying promising regions of the search space. In this section, we discuss how these manual practices can be transformed into algorithmic tools that enable automated early stopping of underperforming runs.
@@
While we can freely choose $b$ during the optimization process, our primary interest lies in the performance at the full budget $b_{max}$. More formally, our goal is (still) to find $\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f_{b_{max}}(\lambda)$, which coincides with the formalization of Chapter 2. Next, we explore commonly used fidelity measures that can be leveraged to accelerate the optimization process.
@@ 5.1.1 Iterations
For iterative machine learning algorithms, such as learning in neural networks, we can interpret the number of iterations as the budget and use the performance at each iteration as an approximation of the final performance [swersky2014freeze, li2017hyperband, jamieson2016non]. In the case of neural networks, we typically consider epochs as iterations rather than individual mini-batch updates to reduce the overhead of validation error computations.
@@ 5.1.2 Training Dataset Subsets
The runtime of most machine learning algorithms scales linearly or even super-linearly with the number of training data points. Figure 5.1 (right) shows the validation error of a support vector machine trained on increasing subsets of the MNIST dataset, with each color representing a different hyperparameter configuration. As expected, the validation error decreases as the training dataset size increases. More interestingly, however, the ranking of hyperparameter configurations remains relatively consistent across dataset subsets [bornschein-icml20]. In the context of deep neural networks, this observation is related to recent research on scaling laws [kaplan-arxiv19, hoffmann-arxiv22], which empirically demonstrate that the training loss of autoregressive transformer models [vaswani_attention_2017] often follows a power-law function with respect to the size of the training dataset.
@@ 5.2.1 Successive Halving
$\mathcal{R}=\{b_{min}\eta^{k}\mid k\in[T]\}.$
We assume, for simplicity, that $b_{max}/b_{min}=\eta^{T}$, where $T\in\mathds{N}$. Here, $\eta\geq 2$ defines the halving constant and determines how aggressively SH prunes configurations.
@@
1:Multi-fidelity response function $f_{b}$
2:Initial budget $b_{min}$, halving constant $\eta\geq 2$, number of rungs $T$
3:Set of $n$ candidate configurations $\mathcal{C}_{0}=\{\lambda_{1},\lambda_{2},\dots\lambda_{n}\}$
4:for $k\leftarrow 1$ to $T$ do
5:  $r_{k}\leftarrow b_{min}\eta^{k}$ $\triangleright$ Define rung level
6:  $\mathcal{F}\leftarrow\{f_{r_{k}}(\lambda)\mid\lambda\in\mathcal{C}_{k-1}\}$ $\triangleright$ Evaluate configurations on the rung
7:  $s\leftarrow\left\lfloor|\mathcal{C}_{k-1}|/\eta\right\rfloor$ $\triangleright$ Number of configurations to keep
8:  $\mathcal{C}_{k}\leftarrow\underset{\lambda\in\mathcal{C}_{k-1}}{\arg\mathrm{top}\text{-}s}\ \mathcal{F}$ $\triangleright$ Keep $s$ best performing configurations
9:end for
Algorithm 2 shows pseudocode for SH. Given a set $\mathcal{C}_{0}$ of $n$ hyperparameter configurations, usually drawn uniformly at random, we evaluate them at the first rung level, i.e., each configuration is evaluated with a budget of $b_{min}$ (Line 2). For convenience, the number of initial configurations is often set to $n=\eta^{T}$, but this can also be arbitrary. Afterwards, SH promotes $\frac{1}{\eta}$ of the configurations, sorted by their validation performance, to the next higher rung level $b_{min}\eta$ (Line 4). These steps are iterated until we reach the highest rung level, $b_{max}$. See Figure 5.2 for a visualization.
While conceptually simple, SH has demonstrated strong performance in the literature and has even outperformed more sophisticated Bayesian optimization techniques that operate only on the full budget, in terms of achieving the same performance in less wall-clock time [jamieson2016non, li2017hyperband]. However, its performance relies on a high correlation between the ranking of hyperparameter configurations across rung levels. In other words, the performance of a configuration at the first rung level should be representative of its performance at the highest rung level. Otherwise, configurations that perform best at $b_{max}$ may be discarded too quickly.
Figure 5.2: Successive halving splits the maximum budget $b_{max}$ into a set of exponentially increasing rungs. Then, in each iteration, successive halving evaluates all configurations on the current rung and promotes the top $1/\eta$ configurations to the next rung. Here, half of the configurations are discarded at every rung, and their learning curve is interrupted. Only two configurations are evaluated for the full budget $b_{max}=100$.
@@ 5.2.2 Hyperband
The minimum budget $b_{min}$ allocated to each configuration is an important meta-parameter of SH. If $b_{min}$ is too small, we may filter out good configurations based on unreliably low fidelities, while if $b_{min}$ is too large, our advantage over standard random search may diminish. Hyperband [li2017hyperband] addresses this issue by running different instances of SH in sequence. The $l$-th instance (or bracket) has rung levels
$\mathcal{R}_{l}=\{b_{min}\eta^{k}\mid k\in\{l,\dots,T\}\},$
so $l=0$ corresponds to SH with $T+1$ rung levels, while $l=T$ corresponds to random search with a single level $b_{max}$. Each bracket has its own number of initial configurations $n_{l}$, which is chosen such that the total budget spent in each bracket is roughly the same. The risk of choosing too small a value for $b_{min}$ is mitigated at a modest extra expense.
@@ 5.2.3 Synchronous Successive Halving
Recall the notation of synchronous versus asynchronous parallelization from Section 2.4, and assume that $W$ workers are available. For synchronous scheduling, we can start $\min(W,M)$ jobs in parallel, where $M$ is the number of free slots remaining in the currently processed rung. Here, WMW-M workers remain idle, which is particularly wasteful for the higher rungs. For example, if $n=\eta^{K}$, the highest rung has only a single slot, so W1W-1 workers will be idle for the maximum budget $b_{max}$. The same issue arises for asynchronous scheduling. A remedy will be discussed in Section 5.2.4.
Hyperband inherits the idle worker issue noted above for synchronous scheduling. However, since brackets are independent of each other, we can work on several of them in parallel, which eliminates idle time for asynchronous scheduling.
@@ 5.2.4 Asynchronous Successive Halving (ASHA)
In ASHA, all available computational resources are utilized immediately, ensuring that workers are never idle. Additionally, superior configurations are quickly promoted toward $b_{max}$ without the need to wait for synchronization points. In practice, ASHA often significantly outperforms synchronous SH or Hyperband in terms of wall-clock time [Klein:20].
@@ 5.3.1 Fabolas
Compared to classical Bayesian optimization (see Chapter 4), we cannot simply optimize standard acquisition functions as described in Section 4.2.3. For example, expected improvement without any further modifications tends to evaluate mostly where $b=b_{max}$, since, as discussed above, here we expect the validation error to be lower than for smaller $b$. As a result, this approach would lead to over-exploitation, primarily evaluating configurations at the highest budget.
@@
$p_{\min}(\lambda^{*}\mid\mathcal{C})=\mathbb{P}\left[\lambda^{*}\in\operatorname*{arg\,min}_{\lambda\in\Lambda}f(\lambda)\right]$
@@ 5.3.2 BO-HB: Bayesian Optimization Hyperband
[…]
@@
While the original implementation of TPE by bergstra2011algorithms used a univariate kernel density estimator, falkner2018bohb proposed using a multivariate kernel density estimator to better capture interactions between hyperparameters. To fit such a model, at least $N_{\min}=m+1$ data points are required, where $m$ is the dimensionality of the search space. Follow up work also considered more sophisticated modelling approaches, such as different GP variants [klein-arxiv20, wistuba-neurips22] or conformalized quantile regression [salinas-icml23].
@@
In this chapter, we discuss a class of randomized methods for global optimization that maintain a population of candidate solution, similar to multi-fidelity methods discussed in the previous chapter. Whereas successive halving and related techniques are rooted in the bandit problem and linked to hedging strategies, the methods we explore here draw inspiration from processes of natural evolution [back1997handbook, simon2013evolutionary]. Evolutionary algorithms date back to the beginning of computer science [neumann1966theory] and have been used to tackle non-smooth and non-convex problems across many domains. Methods in this class typically rely on function evaluations only. Their application to hyperparameter optimization includes methods like CMA-ES [hansen1996adapting], population-based training (PBT) [jaderberg2017population] and particle swarm optimization [lorenzo2017particle], which we review below. Despite the lack of theoretical backing, population-based methods, especially evolutionary algorithms, have enjoyed continued attention in HPO, due to their adaptability to complex search spaces and their strong empirical performances [loshchilov2016cma].
@@ 6.1 Evolutionary Algorithms
Traditionally, evolutionary algorithms (EAs) distinguish between evolutionary strategies (ES) and genetic algorithms (GA). GA refers to a class of methods that deal with binary problems and discrete objective functions while ES is concerned with real-valued, continuous objectives.
The terminology of EA techniques is borrowed from biology: the set of points is called population and configurations are called individuals, chromosomes or genotypes. Iterations are called generations and the objective function is called fitness. Typically, at the end of each generation, the fitness of all individuals is observed, making EAs inherently parallel algorithms — traditional implementations usually requiring synchronization. After observing the fitness of the population, a series of transformations is applied to the individuals. Common types of transformations, which collectively implement explore/exploit strategies in this context, include the following one:
@@
Other strategies like elitism (i.e., keep the best individuals unchanged) and niching (i.e., maintain a diverse population) are often used in conjunction to the explore/exploit strategies outlined above.
We outline this class of methods in Algorithm 3. We can think of individuals and generations as draws from $N$ distributions $\{p_{k}^{j}\}$, where $N$ is the population size. Here, the superscripts $j\in[N]$ index individuals and the subscripts $k\in[T]$ index generations, where $T$ is the number of generations. These distributions implement the strategies described above, and evolve at each generation using global information from the current generation. For instance, elitism can be implemented by setting $p_{k}^{j}=\delta_{\lambda_{k}^{j}}$ for the best performing configurations, where $\delta$ is a Dirac delta distribution. In contrast to random or quasi-random search methods covered in Chapter 3, population-based methods keep in general $N$ distinct distributions rather than a global one that are updated iteratively rather than fixed.
@@
6:   $\lambda_{k}^{j}\sim p^{j}_{k-1}$ $\triangleright$ Draw an individual
7:   Evaluate $f(\lambda_{k}^{j})$ $\triangleright$ Evaluate its fitness (if needed)
@@
9:  Update $p^{j}_{k}$ from $p^{j}_{k-1}$ using $\{(\lambda^{i}_{k},f(\lambda^{i}_{k}))\}_{i\in[N]}$ $\triangleright$ “Evolve”
@@
Various implementations of ES and GA have been proposed in HPO e.g. for tuning hyperparameters of support vector machines [friedrichs2005evolutionary], for neural architecture search [real2019regularized, miikkulainen2024evolving] (see Section 8.3 for a discussion on neural architecture search), parameters of games [kunanusont2017ntuple] and even entire machine learning algorithms [real2020automl]. In the next sections, we review three particular implementations that have received notable attention.